OpenStudy (anonymous):

The point (1,-3) is on a circle with center (-2,1). Write the equation of the circle and graph it as ⊙ C.

OpenStudy (javk):

the eqn of a circle, center (a,b): \[(x-a)^{2}+(y-b)^{2}=r ^{2}\]

OpenStudy (anonymous):

So then (x+2)2+(y-1)=r2?

OpenStudy (javk):

The center of the circle here is (-2,1), so a= -2 and b=1 therefore, plugging in the values \[(x-(-2))^{2}+(y-1)^{2}=r ^{2}\] \[(x+2)^{2}+(y-1)^{2}=r ^{2}\] now you need to figure out what r sqrd is

OpenStudy (anonymous):

Would I use the other coordinates for that?

OpenStudy (javk):

yeah so draw a triangle  and solve using Pythagoras

OpenStudy (javk):

 \[r ^{2}=3^{2}+4^{2}\] \[r ^{2}=25\] \[(x+2)^{2}+(y-1)^{2}=r ^{2}\] replace r sqrd with 25 \[(x+2)^{2}+(y-1)^{2}=25\] and that is yr final ans